OpenStudy (anonymous):

Integration by Parts: Evaluate the Integral of-4x^(2)cos(mx)

zepdrix (zepdrix):

\[\Large\bf \int\limits\color{#DD4747 }{-4x^2} \color{#008353 }{\cos(mx)\;dx}\]We want to destroy those x's. The polynomial part will deteriorate nicely when we take it's derivative, so that will be our u.\[\Large \bf \color{#DD4747}{u=-4x^2}\qquad\qquad\qquad \color{#008353}{dv=cos(mx)\;dx}\] So what will your `du` and `v` be?